# PROFIT: A Specialized Optimizer for Deep Fine Tuning

### v2

@@ Abstract
Fine-tuning pre-trained models has become invaluable in computer vision and robotics. Recent fine-tuning approaches focus on improving efficiency rather than accuracy by using a mixture of smaller learning rates or frozen backbones. To return the spotlight to model accuracy, we present PROFIT (Proximally Restricted Optimizer For Iterative Training), one of the first optimizers specifically designed for incrementally fine-tuning converged models on new tasks or datasets. Unlike traditional optimizers such as SGD or Adam, which make minimal assumptions due to random initialization, PROFIT leverages the structure of a converged model to regularize the optimization process, leading to improved results. By employing a simple temporal gradient orthogonalization process, PROFIT outperforms traditional fine-tuning methods across various tasks: image classification, representation learning, and large-scale motion prediction. Moreover, PROFIT is encapsulated within the optimizer logic, making it easily integrated into any training pipeline with minimal engineering effort. A new class of fine-tuning optimizers like PROFIT can drive advancements as fine-tuning and incremental training become increasingly prevalent, reducing reliance on costly model training from scratch.
@@ 1 Introduction
Fine-tuning pre-trained models has become widely adopted in solving computer vision and robotics problems. As datasets and models both increase in size, having to train a new model for every new application and setting quickly becomes intractable. Imagine, for example, the cost of having to train a new model every time an autonomous vehicle needs to operate in a new city or every time a camera application needs to recognize a new type of object. The shift towards fine-tuning within the deep learning community has also been accelerated by developments in large foundational models that were trained on vast quantities of data, such as CLIP [40] and DINO [5]. Indeed, hints abound that deep learning is steadily inching towards a new paradigm where only simple models are trained from scratch.
@@
At the same time, fine-tuning a model comes with its own set of challenges. For one, it is known that models tend to readily forget old information when fine-tuned on new information, in a process known as Catastrophic Forgetting [16]. Various mitigation methods have been proposed [30], but require additional data engineering and modifications of the model architecture. The common practice within fine-tuning still largely relies on training on new tasks/data with a smaller learning rate or with a frozen backbone (or both). Parameter-efficient fine-tuning methods such as LoRA [18] introduce learnable adapter modules for transfer learning, but serves mainly to improve fine-tuning efficiency rather than accuracy [2]. We aim to specifically improve the accuracy of fine-tuning in the present work, while keeping the process efficient by not introducing any additional parameters.
@@
One ubiquitous element within deep learning training that is also almost always implemented in a modular fashion is the optimizer. The majority of models set the optimizer to some noncontroversial choice (such as Adam  [24], AdamW [34] or Momentum [47]). Yet, all current optimizers are designed with training-from-scratch in mind, which forces their design to assume minimal structure about the problem setting. In contrast, the fine-tuning setting usually starts from a well-trained, well-converged model that we already trust. So we ask: how do we design an optimizer especially for the case where we start from a converged model from a similar domain? Given the ubiquity and modularity of optimizers within training pipelines, such an optimizer would be immediately applicable and easy to implement within any fine-tuning setting.
@@
Works such as Learning Without Forgetting (LWF) [30] have proposed mitigating catastrophic forgetting by enforcing proximity to an old state of the model, but requires additional data pipelining and model snapshots that serve as additional supervision to keep the model anchored to its old “good” state. Instead of anchoring a model across tasks, we propose a different but equally valid anchoring across time. For each iteration of the optimizer, we can take a step away from equilibrium and balance further steps away from equilibrium with the model’s desire to return to equilibrium using standard methods from the gradient-based multitask learning literature. The result is a system that mimics data-driven anchor methods such as [30] without actually needing data processing and replaces the rigid static anchors of other methods with a dynamically updated flexible one.
Figure 1: Schematic of a gradient update for PROFIT (Proximally Restricted Optimizer For Iterative Training). Unlike standard optimizers, which take successive steps through parameter space, PROFIT keeps track of a “good” model state and collects reference information of how the system tries to move away from the “good” state (denoted by the superscript “r” - for “reference” - in the figure). The functionality of the “good” state is then restored by translating and orthogonalizing gradient updates with respect to that state. These processes allow PROFIT to take heavily regularized update steps and produce superior performance when in challenging fine-tuning settings. The visualization in the bottom row is discussed in Section 4.1.
@@
Specifically, a model converged in some state $\mathbf{x}_{0}$ will proceed along the states $\mathbf{x}_{1},\ldots\mathbf{x}_{t}$ with a standard optimizer. A further update would send the system to $\mathbf{x}_{t+1}$. However, because $\mathbf{x}_{0}$ was a “good state,” the model also would benefit by returning to $\mathbf{x}_{0}$. We thus have two potentially conflicting gradient directions ($\mathbf{\Delta}:=\mathbf{x}_{t}\rightarrow\mathbf{x}_{0}$ and $\mathbf{g}:=\mathbf{x}_{t}\rightarrow\mathbf{x}_{t+1}$), which is a classic multitask learning problem. We borrow from [55] and assign $\mathbf{g}\mapsto\mathbf{g}\perp\mathbf{\Delta}$, where $\perp$ is the orthogonalization operator. We restore the model to state $\mathbf{x}_{0}$ (“translate” operation on the top right of Figure 1) and take a step in the orthogonalized direction $\mathbf{g}$.
@@
The operation of PROFIT (Proximally Restricted Optimizer For Iterative Training) relies on compatibility between the fine-tuning and baseline settings, as it aims to keep the model state close to the baseline dynamically. PROFIT excels at adapting a model to a nearby distribution, making this compatibility a crucial aspect of our approach. This excludes some cases often referred to as fine-tuning (e.g., pre-training on ImageNet, fine-tuning on unrelated datasets), we distinguish these pre-training scenarios from practical fine-tuning under compatibility constraints. We refer to this approach as “proximal fine-tuning”, where the fine-tuning dataset is from a similar distribution to pre-training. Although this constraint may seem limiting, “proximal fine-tuning” is prevalent in machine learning applications. For instance, a self-driving car’s trajectory prediction network may need to train a prediction head for a new scooter type using the same sensors.
We name our method PROFIT, which stands for PROximal FIne Tuning, and a schematic of the process is shown in Figure 1. To the best of our knowledge, PROFIT is one of the first optimizers that is specifically designed to work in the fine-tuning setting. PROFIT provides significant performance boosts in those cases, as we will demonstrate.
Our main contributions are as follows:
@@
  • We present PROFIT, an optimizer designed specifically for fine-tuning a good baseline model, and which can be easily dropped into any deep learning training framework.
  • We prove that PROFIT allows us to train as if we still had access to the old data/setting in an unsupervised way.
@@
  • We experimentally show that PROFIT beats standard fine-tuning methods in a large number of settings, from a low-dimensional toy example to image classification to large-scale motion prediction for autonomous driving.
@@ 2 Related work
For detailed background context in multi-task learning (MTL), we refer the reader to [59]. MTL [59] is an optimization problem where we train a model on multiple tasks to take advantage of the structure of shared neural networks, thus improving generalization. One direction is to use gradient descent method to optimize joint multi-task learning problem. Our setting shares some similarities with that. The difference is that our problem is considered as temporal multi-task learning. Ozan [44] formulates the MTL problem as a multi-objective optimization problem and learns the loss weights that change dynamically. GradNorm [10] attempts to normalize gradients to balance learning of multiple tasks. PCGrad [55] suggests that to mitigate gradient direction conflicts, we should project a task’s gradient onto the normal plane of the gradient of any other tasks where a gradient conflict is present. Our proposed technique is similar to PCGrad, but modifies the core designs specifically for the “proximal fine-tuning” setting.
Adapters
An adapter introduces learnable parameters to help transfer learning. The number of trainable parameters in the adapter is much smaller than the model itself, which makes them appealing in practice. Several strategies have been proposed such as visual prompt tuning [21], side adapters [57], bias tuning [4], and residual adapters [41] for efficient transfer learning. LoRA [18] updates the projection matrices in an attention module by decomposing them to a low-rank matrix. In contrast to our method, adapters require additional parameters and may also assume a specific class of model architectures for their success (e.g: transformers).
@@
Fine-tuning on particular datasets [27, 9] is a common technique in the era of deep learning. Practitioners often use standard optimizers such as Stochastic Gradient Descent (SGD) [3], Adam [24], and AdamW [34]. Recent architectures such as ViTs [12, 5, 40] or ConvNeXts [32] use AdamW for fine-tuning, while it is also common to use SGD for fine-tuning models like ResNets [17, 26] due to the optimizer’s efficiency. However, SGD and AdamW do not assume that we want to stay close to our model’s start state, and thus lead to models that tend to compromise performance on old data when fine-tuning on new data, also known as catastrophic forgetting  [37]. Our method serves as a regularization approach to bridge the gap in existing optimizers to mitigate this forgetting issue.
@@
To address this issue, various approaches have been developed, such as regularization-based methods [25, 8, 22, 48, 19], where the goal is to keep learned information of the past tasks during continual learning. In particular, Mirzadeh et al. [19] demonstrated a link between continual learning and multi-task learning, and Farajtabar et al. [14] proposed projecting gradients from new tasks onto the subspace of prior task gradients. Learning Without Forgetting (LWF) [30] proposes storing the old model response on new tasks/data and supervises the new model on these responses using a distillation loss to prevent catastrophic forgetting [36]. Our work draws inspiration from LWF’s key idea of storing a pretrained model’s response, but we aim to avoid the storage of old data/checkpoint/statistics. Instead, we use the network’s response at the initial state of each iteration. Two other directions to address catastrophic forgetting are rehearsal-based methods [42, 6, 33, 7, 43] that directly make use of the old data source, and architecture-based methods that minimize intertask interference through new architectures [35, 45, 29, 52, 53]. These methods add substantial infrastructural overhead, while our approach is attractive in its simplicity of implementation. Since our method is an optimizer by design and does not require any additional storage (unlike the continual learning paradigm), we do not benchmark against continual learning methods.
@@ 3.1 The PROFIT Optimizer
We describe the complete algorithm of PROFIT in Alg 1. Our method is easy-to-understand and modular, which can be encapsulated within the definition of an optimizer step. Next, we provide a discussion of the hyperparameters and the details of the implementation, and conclude with our assumptions and the limitations of PROFIT.
@@
6:     $\mathbf{\theta}_{\text{ref}}\leftarrow\mathbf{\theta}$triangleright Save the model state.
@@
16:         $\mathbf{g}\leftarrow\mathbf{g}\perp\Delta$triangleright $\mathbf{a}\perp\mathbf{b}$ denotes orthogonal projection of $\mathbf{a}$ onto $\mathbf{b}$
@@
19:     $\theta\leftarrow\mathbf{O}(\theta,\mathbf{g})$triangleright Take step with main optimizer.
@@
At a high level, an implementation of PROFIT involves defining an optimizer wrapper which takes as argument two separate standard optimizers, $\mathbf{O}$ and $\mathbf{O}^{(\text{ref})}$ as inputs. The latter “reference” optimizer perturbs the system from equilibrium, while the former “main” optimizer uses this perturbation to make a final update. The entire logic of PROFIT can be encapsulated within the logic of this optimizer wrapper class, making the method very portable, modular, and invariant to the choice of “main” optimizer.
@@
Given a model $\mathcal{M}$ with weights $\theta$ which has been trained on data $\mathbf{X}$ corresponding to an old task, and a data source $\mathbf{X^{\prime}}$ from a similar domain corresponding to a new task (with a task loss $L$), we aim to fine-tune $\mathcal{M}$ to work well on both the old and new tasks. Note that we make a strict assumption about having a converged model $M$ available as an input to PROFIT; any random or untrained weights in the model can lead to poor performance. Additionally, our model also requires two instantiated optimizers, a standard optimizer $\mathbf{O}$ and a reference optimizer $\mathbf{O}^{(\text{ref})}$. The user can tune $\mathbf{O}$ as per their needs, while we recommend using SGD for $\mathbf{O}^{(\text{ref})}$.
@@
A step through our optimizer works as follows. First, we store the current state of $\mathcal{M}$ by saving $\theta$ into $\theta_{\text{ref}}$. Next, we draw $n_{\text{ref}}$ batches from $\mathbf{X^{\prime}}$, one at a time, and iteratively minimize $L$ with the reference optimizer $\mathbf{O^{(\text{ref})}}$. We have now perturbed the system from equilibrium and must decide the best way to restore the said equilibrium.
@@
To do so, we first calculate $\Delta:=\theta-\theta_{\text{ref}}$, the total displacement of our original position after $n_{\text{ref}}$ steps of the optimizer $\mathbf{O^{(\text{ref})}}$. We reason that if the stored equilibrium state corresponds to a good critical point of the original model, then $\Delta$ corresponds to the benign gradient direction that will restore the original critical point. We thus have two potentially conflicting gradient updates: the one corresponding to $\Delta$, and the one corresponding to the next queried update by $\mathbf{O^{(\text{ref})}}$, which we call $\mathbf{g}$. We need to decide on how to take a single gradient step that is consistent with both options.
@@
We then borrow an idea from PCGrad [55], which reconciled conflicting gradients by orthogonally projecting them onto each other in a pairwise fashion. Crucially, we choose to only project $\mathbf{g}$ onto $\Delta$, and not the other way around, because $\Delta$ represents a gradient on the old dataset which may no longer be accessible and, therefore, must be treated with more care. Therefore, we end with the two gradient updates $\mathbf{g}$ and $\mathbf{g}\perp\Delta$, and we take both steps by first restoring $\theta\mapsto\theta_{\text{ref}}$ and then allowing $\mathbf{O}$ to take a step in the $\mathbf{g}\perp\Delta$ direction. This process is repeated until training is complete.
@@
Interestingly, this formulation allows us to view fine-tuning as temporal multi-task learning, with the two tasks being: (1) pretraining ($\Delta$), and (2) fine-tuning ($\mathbf{g}$). To the best of our knowledge, this is the first time it has been viewed through this lens, and this insight may pave avenues for future research on fine-tuning.
3.2 Hyperparameter Discussion and Tradeoffs
PROFIT introduces three main hyperparameters for fine-tuning: $n_{\text{ref}}$, $\mathbf{O}^{\text{(ref)}}$, and $\lambda_{\text{ref}}$. As previously discussed, $n_{\text{ref}}$ controls the degree of exploration for the reference optimizer, while $\lambda_{\text{ref}}$ controls the step size at each iteration of the reference step. The choice of $\mathbf{O}$ can be set by the practitioner based on their needs, but $\mathbf{O}^{(ref)}$ should generally be set to standard SGD, as the gradient calculations that drive PROFIT are cleanest when the reference updates are simple.
The primary concern for practitioners using PROFIT is the cost of setting $n_{\text{ref}}$, as it requires $n_{\text{ref}}$ additional optimization steps per training step. In practice, the practitioner is encouraged to start with $n_{\text{ref}}=1$, which will often work well, and only to increase the value if they observe the performance staying flat. However, we note two reasons why the potentially increased compute of PROFIT is not a large issue: (1) we find that PROFIT generally converges quicker, possibly due to the positive regularization effects of the method, and (2) fine-tuning is generally run for much fewer steps than training from scratch, so the effects of this additional training time are often still very tractable. For all experiments, we use $n_{\text{ref}}=1$.
Mathematically, $\lambda_{\text{ref}}$ encodes how quickly we travel from equilibrium to study the general shape of the loss surface, so that we can make an informed decision on where to go next. In practice, we find that there is often a sweet spot somewhere between $\lambda_{main}/10$ and $\lambda_{main}/10000$, but the exact value is heavily dependent on the loss surface for the specific problem. It may in fact turn out that the optimal value of $\lambda_{\text{ref}}$ tells us something about the fundamental properties of that particular setting’s loss surface, but such analysis falls outside the scope of the present work.

### v3

@@ Abstract
The fine-tuning of pre-trained models has become ubiquitous in generative AI, computer vision, and robotics. Although much attention has been paid to improving the efficiency of fine-tuning models, there has been less scholarship around fine-tuning specifically for improved model performance. To remedy this gap, we present PROFIT, one of the first optimizers designed to incrementally fine-tune converged models on new tasks and/or datasets. Unlike traditional optimizers such as SGD or Adam, which make minimal assumptions due to random initializations, PROFIT takes the properties of a converged model into account explicitly to regularize the optimization process. Employing a temporal gradient-orthogonalization process, PROFIT outperforms fine-tuning methods in various tasks, from image classification to multimodal language model training to large-scale motion prediction. Moreover, PROFIT is encapsulated as a modular optimizer, which makes it easy to integrate directly into any training pipeline with minimal engineering effort.
@@ 1 Introduction
Fine-tuning pre-trained models has become widely adopted in solving computer vision and robotics problems as well as in modern generative AI settings. As datasets and models increase in size, having to train a new model for every new application and setting quickly becomes intractable. Imagine, for example, the cost of having to train a new model every time an autonomous vehicle needs to operate in a new city, or every time a camera application needs to recognize a new type of object, or every time a custom LLM needs to update its knowledge cutoff. The shift towards fine-tuning within the deep learning community has also been accelerated by developments in large foundational models that were trained on vast quantities of data, such as CLIP Radford et al. (2021), DINO Caron et al. (2021), and open-source LLMs Touvron et al. (2023a). Indeed, deep learning is steadily inching towards a new paradigm where very few practitioners are training models from scratch.
@@
At the same time, fine-tuning a model comes with its own set of challenges. For one, models are known to readily forget old information when fine-tuned with new information, in a process known as catastrophic forgetting Goodfellow et al. (2013). Various mitigation methods have been proposed Li and Hoiem (2017), but require data engineering and modifications to the model architecture. Common practice within fine-tuning still largely relies on training on new tasks/data with a smaller learning rate or with a frozen backbone (or both). Parameter-efficient fine-tuning methods such as LoRA Hu et al. (2022) introduce learnable adapters for transfer learning, but serve to improve fine-tuning efficiency rather than accuracy (Biderman et al. (2024)). We aim to improve the accuracy of fine-tuning while keeping the process efficient by not introducing any additional parameters.
@@
A ubiquitous object within deep learning that is modular and abstracted away from the practitioner is the optimizer. Many AI models set the optimizer to popular standards like Adam  Kingma and Ba (2015), AdamW Loshchilov and Hutter (2019), or Momentum Sutskever et al. (2013). However, all current optimizers are designed for training from scratch, so they make minimal assumptions about the problem setting and initial model state. In contrast, the fine-tuning setting usually starts from a well-trained, well-converged model that already performs well on some set of meaningful data. So we ask: how do we design an optimizer specifically to start from a converged model from a similar domain? Given the ubiquity and modularity of optimizers within training pipelines, such an optimizer would be immediately applicable and easy to implement within any fine-tuning setting.
@@
Works such as Learning Without Forgetting (LWF) Li and Hoiem (2017) have proposed mitigating catastrophic forgetting by enforcing proximity to an old state of the model, but require additional data pipelining and model snapshots that serve as additional supervision to keep the model anchored to its old “good” state. Instead of anchoring a model across tasks, we propose a different but equally valid anchoring across time rather than across tasks. For each iteration of the optimizer, we take a step away from equilibrium and balance further steps away from equilibrium with the model’s desire to return to equilibrium using techniques from the gradient-based multi-task learning literature. The result is a system that mimics data-driven anchor methods such as Li and Hoiem (2017) without incurring any additional data processing, and which foregoes the rigid static anchors of more classic methods in favor of a dynamically updated flexible one.
Figure 1: Schematic of PROFIT. Standard fine-tuning (middle) takes successive steps away from a good starting state $\theta_{0}$. PROFIT (right): (1) take $n_{\text{ref}}$ small reference steps with $O^{(\mathrm{ref})}$ to obtain a displaced state $\theta_{1}$; (2) compute the displacement $\Delta=\theta_{1}-\theta_{0};$ (3) orthogonalize the new-batch gradient $g:=\theta_{2}-\theta_{1}$ to $\Delta$; (4) restore $\theta\leftarrow\theta_{0}$ and then apply the main optimizer $O$ along the orthogonalized direction. Here, r denotes ‘reference’. See Alg. 1 for details.
@@
Specifically, a model converged in some state $\mathbf{\theta}_{0}$ will proceed along the states $\mathbf{\theta}_{1},\ldots\mathbf{\theta}_{t}$ with a standard optimizer. A further update would send the system to $\mathbf{\theta}_{t+1}$. However, because $\mathbf{\theta}_{0}$ was a “good state,” the model would also benefit by returning to $\mathbf{\theta}_{0}$ (Sec. 3). We thus have two potentially conflicting gradient directions ($\mathbf{\Delta}:=\mathbf{\theta}_{t}\rightarrow\mathbf{\theta}_{0}$ and $\mathbf{g}:=\mathbf{\theta}_{t}\rightarrow\mathbf{\theta}_{t+1}$), which is a classic multitask learning problem. We borrow from Yu et al. (2020) and assign $\mathbf{g}\mapsto\mathbf{g}\perp\mathbf{\Delta}$, where $\perp$ is the orthogonalization operator. We restore the model to state $\mathbf{\theta}_{0}$ (“translate” operation on the top right of Figure 1) and take a step in the orthogonalized direction $\mathbf{g}$.
@@
The operation of PROFIT relies on some non-negligible distributional overlap between the fine-tuning and baseline settings, as it aims to dynamically keep the model state close to the baseline state. This requirement precludes some settings that are common within fine-tuning like pre-training and then fine-tuning on completely orthogonal datasets. We refer to our approach with this additional constraint as “proximal fine-tuning”, where the fine-tuning dataset is of a similar distribution to that of the pre-training dataset. Although this constraint may seem limiting, “proximal fine-tuning” is prevalent in machine learning applications. For example, a self-driving car’s trajectory prediction network may need to train a prediction head for a new scooter type using the same sensor, while maintaining performance on old tasks. We will also later show how to overcome this constraint even in non-proximal settings by introducing a warmup phase.
PROFIT (PROximal FIne Tuning) is shown schematically in Fig. 1. To the best of our knowledge, PROFIT is among the first optimizers explicitly designed for fine-tuning.
Our main contributions are as follows.
@@
  • We introduce PROFIT, an optimizer for fine-tuning converged models, easily integrated into any deep learning framework.
  • We show that PROFIT allows unsupervised training as if the original data were available.
@@
  • We show that PROFIT outperforms standard fine-tuning methods on various tasks, from image classification to VLM fine-tuning to large-scale motion prediction for autonomous driving.
@@ 2 Related work
For detailed background context in multi-task learning (MTL), we refer the reader to Zhang and Yang (2021). MTL Zhang and Yang (2021) is an optimization problem in which we train a model on multiple tasks simultaneously to take advantage of the structure of shared neural networks, thus improving generalization and efficiency. One direction is to use gradient descent methods to optimize the joint multi-task learning problem. Our setting shares some similarities with MTL, with the key difference that problem is a temporal multi-task learning. Ozan Sener and Koltun (2018) formulates the MTL problem as a multi-objective optimization problem and learns the loss weights that change dynamically. GradNorm Chen et al. (2018) attempts to normalize gradients to balance learning of multiple tasks. PCGrad Yu et al. (2020) suggests that to mitigate gradient direction conflicts, we should project a task’s gradient onto the normal plane of the gradient of any other task where a gradient conflict is present. Our proposed technique is similar to PCGrad, but modifies the core designs specifically for the “proximal fine-tuning” setting.
Parameter-efficient fine-tuning
Parameter-efficient fine-tuning reduces computational and memory requirements by updating fewer parameters, which simplifies the adaptation of large models to new tasks. An adapter Houlsby et al. (2019) introduces light-weight learnable parameters to help transfer learning. The trainable parameters in the adapter are much smaller than in the model, which makes them attractive in practice. Several strategies have been proposed, such as visual prompt tuning Jia et al. (2022), side adapters Zhang et al. (2020), bias tuning Cai et al. (2020), and residual adapters Rebuffi et al. (2017a) for efficient learning. LoRA Hu et al. (2022) decomposes the matrices of an attention module into a low-rank matrix. In contrast to our method, adapters require additional parameters and may also assume a specific class of model architectures for their success (e.g., transformers).
@@
Fine-tuning on particular datasets Kornblith et al. (2019); Chen et al. (2020) is a common technique in the era of deep learning. Practitioners often use standard optimizers such as Stochastic Gradient Descent (SGD) Bottou (2010), Adam Kingma and Ba (2015), and AdamW Loshchilov and Hutter (2019). Recent architectures such as ViTs Dosovitskiy et al. (2021); Caron et al. (2021); Radford et al. (2021) or ConvNeXts Liu et al. (2022) use AdamW for fine-tuning, while it is also common to use SGD for fine-tuning models like ResNets He et al. (2016); Kolesnikov et al. (2020) due to the optimizer’s efficiency. However, SGD and AdamW do not assume that we want to stay close to our model’s start state, and thus lead to models that tend to compromise performance on old data when fine-tuning on new data, also known as catastrophic forgetting  McCloskey and Cohen (1989). Our method serves as a regularization approach to bridge the gap in existing optimizers to mitigate this forgetting issue.
@@
Various approaches have been developed Kirkpatrick et al. (2017); Chaudhry et al. (2020); Jung et al. (2020); Titsias et al. (2020); Mirzadeh et al. (2021), in which the goal is to keep the information learned from the past tasks during continual learning. Learning Without Forgetting (LWF) Li and Hoiem (2017) stores the response of the old model on new tasks/data and supervises the new model on these responses using distillation to prevent catastrophic forgetting Masana et al. (2022). Our work draws inspiration from LWF’s key idea of storing a pre-trained model’s response, but we aim to avoid the storage of old data/checkpoint/statistics. Instead, we use the network’s response at the initial state of each iteration. Other directions for addressing catastrophic forgetting are rehearsal-based methods Rebuffi et al. (2017b); Chaudhry et al. (2019a); Lopez-Paz and Ranzato (2017); Chaudhry et al. (2019b); Saha et al. (2022) that directly make use of the old data source, and architecture-based methods that minimize inter-task interference through new architectures Mallya and Lazebnik (2018); Serrà et al. (2018); Li et al. (2019); Wortsman et al. (2020); Wu et al. (2019). These methods add substantial infrastructural overhead, while our approach is attractive in its simplicity. Since PROFIT does not require replay buffer or architectural changes, we primarily benchmark against optimizers and fine-tuning techniques that practitioners would otherwise use.
@@ 3.1 The PROFIT Optimizer
6:  $\mathbf{\theta}_{\text{ref}}\leftarrow\mathbf{\theta}$ $\triangleright$ Save the model state.
@@
16:   $\mathbf{g}\leftarrow\mathbf{g}\perp\Delta$ $\triangleright$ $\mathbf{a}\perp\mathbf{b}$ denotes orthogonalizing $\mathbf{a}$ with respect to $\mathbf{b}$
@@
19:  $\theta\leftarrow\mathbf{O}(\theta,\mathbf{g})$ $\triangleright$ Take step with main optimizer.
@@
Implementing PROFIT (Algorithm 1) involves defining an optimizer wrapper that takes two standard optimizers $\mathbf{O}$ and $\mathbf{O}^{(\text{ref})}$ as inputs. The latter “reference” optimizer perturbs the system from equilibrium, while the former “main” optimizer uses this perturbation to make the final update. The entire logic of PROFIT can be encapsulated within the logic of this optimizer wrapper class, making the method very portable, modular, and invariant to the choice of “main” optimizer.
@@
Given a model $\mathcal{M}$ with weights $\theta$ trained on data $\mathbf{X}$ corresponding to an old task, and a data source $\mathbf{X^{\prime}}$ from a similar domain corresponding to a new task (with a task loss $L$), our objective is to fine-tune $\mathcal{M}$ to work well on both old and new tasks. We make a strict assumption about having a converged model $M$ available as input to PROFIT; untrained weights will lead to poor performance. Furthermore, our approach also requires two optimizers, a standard optimizer $\mathbf{O}$ and a reference optimizer $\mathbf{O}^{(\text{ref})}$. The user can tune $\mathbf{O}$ as per their needs, while we recommend using SGD for $\mathbf{O}^{(\text{ref})}$.
@@
A step through our optimizer works as follows. First, we store the current state of $\mathcal{M}$ by saving $\theta$ in $\theta_{\text{ref}}$. Next, we draw $n_{\text{ref}}$ batches from $\mathbf{X^{\prime}}$, one at a time, and iteratively minimize $L$ with the reference optimizer $\mathbf{O^{(\text{ref})}}$. We have now perturbed the system from equilibrium and must decide the best way to restore the said equilibrium.
@@
To do so, we first calculate $\Delta:=\theta-\theta_{\text{ref}}$, the displacement vector after $n_{\text{ref}}$ steps of the optimizer $\mathbf{O^{(\text{ref})}}$. We reason that if the stored equilibrium state corresponds to a good critical point of the original model, then $\Delta$ corresponds to the benign gradient direction that will restore the original critical point, as gradient descent would send us in the $\Delta:=\theta_{\text{ref}}-\theta$ direction. We thus have two potentially conflicting gradient updates: the one corresponding to $\Delta$, and the one corresponding to the next queried update by $\mathbf{O^{(\text{ref})}}$, which we call $\mathbf{g}$. We need to decide on how to take a single gradient step that is consistent with both options.
@@
We then borrow an idea from PCGrad Yu et al. (2020), which reconciled conflicting gradients by orthogonally projecting them onto each other in a pairwise fashion. Crucially, we choose to project only $\mathbf{g}$ onto $\Delta$, and not the other way around, because $\Delta$ represents a gradient towards the old dataset that may no longer be accessible and therefore must be treated with more care. We end with the two gradient updates $\mathbf{g}$ and $\mathbf{g}\perp\Delta$, and take both steps by first restoring $\theta\mapsto\theta_{\text{ref}}$ and then allowing $\mathbf{O}$ to take a step in the $\mathbf{g}\perp\Delta$ direction. This process is repeated until training is complete.
@@
This formulation allows us to view fine-tuning as temporal multi-task learning, with the two tasks being: (1) pretraining ($\Delta$), and (2) fine-tuning ($\mathbf{g}$). To the best of our knowledge, this is the first time it has been viewed through this lens, and this insight may pave avenues for future research.